OpenStudy (anonymous):

CAN SOMEONE PLLLZZZZ HELP!!!!!!!!!!!! divide (2x/5y)/(x^2/25)

OpenStudy (whpalmer4):

\[\large{\frac{\frac{2x}{5y}}{\frac{x^2}{25}}}\] You divide fractions by inverting the denominator fraction and multiplying it with the numerator fraction. So this one becomes \[\frac{2x}{5y}*\frac{25}{x^2} =\]

OpenStudy (anonymous):

everyone is saying (10)y/x but thats not the right anwer

OpenStudy (whpalmer4):

And you can see that not only did @fernmndez violate the http://openstudy.com/code-of-conduct, they gave you an incorrect answer...

OpenStudy (anonymous):

would it become 50x/5yx^2

OpenStudy (whpalmer4):

\[\frac{2x}{5y}*\frac{25}{x^2} = \frac{2*x*5*5}{5*y*x*x} = \frac{2*\cancel{x}*5*\cancel{5}}{\cancel{5}*y*x*\cancel{x}} = \]

OpenStudy (anonymous):

ok so it would be 10/yx

OpenStudy (whpalmer4):

@ineedhelp72 Some would accept that, others would not. I claim that you should either write 10/(yx) or learn how to typeset it as \[\frac{10}{yx}\]because it could also be interpreted as \[\frac{10}{y}*x\] otherwise.

OpenStudy (whpalmer4):

No point in getting the right answer if you can't clearly communicate it!